# A high-order recombination algorithm for weak approximation of stochastic differential equations

### v1

@@ 2.3. N–V method
We introduce the N–V method from [24], which is a second-order discretisation. We denote by $(G_{1}\circ G_{2})g(x)$ the composition $(G_{1}(G_{2}g))(x)$ of $G_{1}$ and $G_{2}:C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})\rightarrow C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$ hereinafter.
@@
Let $\eta_{1},\eta_{2},\dots,\eta_{d}$ be independent standard normal random variables and
@@
Then $Q^{(\rm{NV})}_{(s)}$ are 5-similar operators and the N–V method $Q^{(\rm{NV})}_{(T/n)}\circ Q^{(\rm{NV})}_{(T/n)}\circ\cdots\circ Q^{(\rm{NV})}_{(T/n)}$ is a second order discretisation method.
Replacing the standard normal random variables ${\eta}$ with the discrete random variables $\hat{\eta}^{(\text{$5$th})}$, we define the discrete N–V method as in  [22]. Let $\hat{\eta}^{(\text{$5$th})}$ be a discrete random variables on $(\hat{\Omega},\hat{\mathcal{F}},\hat{P})$ defined as
$\hat{P}\left(\hat{\eta}^{(\text{$5$th})}=\pm\sqrt{3}\right)=\frac{1}{6},\;\;\hat{P}\left(\hat{\eta}^{(\text{$5$th})}=0\right)=\frac{2}{3}.$ (22)
We note that the moments of $\hat{\eta}^{(\text{$5$th})}$ up to the $5$th order are equal to those of the standard normal random variables. As is subsection 1.2.2, we denote a probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}\left(t_{i}\right)$ as $\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}$, where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1\dots,m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}$ in $\mathbb{R}^{N}$ and weights $w_{j}^{(i)}$ assigned to them. Let $\{\hat{\eta}^{(\text{$5$th})}_{i,l}\}_{\begin{subarray}{c}i=1,\dots,n\\ l=1,\dots,d\end{subarray}}$ be a set of independent and identically distributed discrete random variables following the same distribution as $\hat{\eta}^{(\text{$5$th})}$. Let $\phi$ be a function from {±$\{\pm\sqrt{3},0\}^{d}$ to $[0,1]$ defined by
$\phi(\eta)=\left(\frac{1}{6}\right)^{\operatorname{card}\{i\;|\;\eta_{i}=\pm\sqrt{3}\}}\left(\frac{2}{3}\right)^{\operatorname{card}\{i\;|\;\eta_{i}=0\}}$
for $\eta=\left(\eta_{1},\dots,\eta_{d}\right)\in\{\pm\sqrt{3},0\}^{d}$. We also define functions $y^{(+)},y^{(-)}$ from $\mathbb{R}^{N}\times\{\pm\sqrt{3},0\}^{d}\times\mathbb{R}_{>0}$ to $\mathbb{R}^{N}$ as
$\begin{split}y^{(+)}\left(x,\eta,s\right)&=\exp\left(sV_{0}\right)\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\cdots\exp\left(\sqrt{s}\eta_{d}V_{d}\right)\left(x\right),\\ y^{(-)}\left(x,\eta,s\right)&=\exp\left(\sqrt{s}\eta_{d}V_{d}\right)\cdots\exp\left(\sqrt{s}\eta_{1}V_{1}\right)\exp\left(sV_{0}\right)\left(x\right).\end{split}$
Then, the operator $\hat{Q}_{(s_{i})}^{({\rm NV},\Delta)}$ between discrete measures can be defined as
@@
In this way, a concrete example of $\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{i})}$ used in this paper, as discussed in subsection 1.2.2, is constructed. We apply the recombination algorithm described in subsection 1.3.2 to this operator and obtain $\mu_{T}^{\left({\rm NV},\Delta,{\rm Rec}\right)}$.
Similarly, for an $m$-moment similar operator $Q_{(s)}$, we define $\hat{Q}_{(s_{i})}(\mu)$ by adapting discrete the random variables $\hat{\eta}^{(m\text{th})}$ whose moments up to the $m$th order are matched with those of the standard normal random variables. In this paper, we consider a $p$th order discretisation algorithm defined from $p$-moment similar operator and its recombination measure.
@@
In [24], the authors propose to implement the N–V method using the Bernoulli random variables in addition to the standard normal random variables $\eta_{1},\eta_{2},\dots,\eta_{d}$, in order to halve the number of ODEs to be solved. In contrast, the present paper employs the N-V method without utilising Bernoulli random variables, instead calculating the expectation as shown in equation (21). This approach is motivated by the necessity to minimise the number of support points of the discrete N–V method in order to regulate the overall calculation cost, as noted in Remark 3.
@@
The N–V method has the following advantages compared to other high-order discretisation methods such as [20, 26]. As stated in the original paper [24], the N–V method is often more efficiently implementable, because analytical solutions for ODEs are available in some cases such as the Heston model [10]. Furthermore, the Feller condition [5] ensures the positivity of a process discretised by the N–V method.
Moreover, when we implement the high-order discretisation methods with discrete random variables, the number of random variables involved in the N–V method is parsimonious in the sense that only $d$ random numbers are required to approximate an SDE driven by the $d$-dimensional Brownian motion. The numbers of random variables and supports of the measures are summarized in Table 1.
@@
Table 1. Number of supports of $\hat{X}^{(\text{method})}_{s,x}$, $\sharp\Delta=n$
@@ 2.4. High-order recombination: approximation theorem
In this subsection, we introduce the approximation theorem established by [17]
@@
The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the uniform Hörmandar (UH) condition, if there exists $\ell\in\mathbb{N}$ such that
@@
Theorem 4 ([17] Theorem 19).
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th order discretisation and $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let $\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=0,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ be a set of recombination measures and patch divisions of their supports. Then,
$\left\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right\|_{\infty}\\ \leq\left(C_{1}(T)\left(s_{\sharp\Delta}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{j=m}^{m+1}\frac{s_{i}^{(j+1)/2}}{(T-t_{i})^{j/2}}\right)+C_{2}(T)\sum_{i=2}^{\sharp\Delta-1}\frac{u_{i}^{m+1}}{(T-t_{i})^{m\ell}}\right)\|f\|_{\rm{Lip}}$ (24)
holds where $u_{i}=\max\left.\left\{\operatorname{rad}\left(U_{k}^{(i)}\right)\;\right|\;k=1,\dots,l_{i}\right\}$ and $s_{i}=t_{i}-t_{i-1}$.
Example 1 (Patch condition ([17] Examples 21)).
@@
$t_{i}=T\left(1-\left(1-\frac{j}{n}\right)^{\gamma}\right)$
@@
$u_{i}=\left(\frac{s_{i}^{m+1}}{(T-t_{i})^{m-r\ell}}\right)^{1/(2(l+1))},$
@@
In other word, the recombination do not harm the order of discretisation error.
@@ 3. Refined error estimation and patch division algorithms
In this section, we present the main results of this paper. First, we refine Theorem 4 and Example 1 by extending the argument in [17] to obtain Theorem 5 and an example of a sufficient condition (Example 2) for the patch division to satisfy the required recombination error.
@@ 3.1. Refined error estimation
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th order discretisation and $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let $\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=0,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ be a set of recombination measures and patches divisions of their supports. Then,
$\begin{split}&\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\|_{\infty}\\ &\leq\Biggl{(}C_{1}(T)\left(s_{\sharp\Delta}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{q=m+1}^{K}\frac{s_{i}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\\ &\qquad\qquad\qquad+C_{2}(T)\sum_{i=1}^{\sharp\Delta-1}\frac{\sum_{k=1}^{l_{i}}\left(\operatorname{rad}\left(U_{k}^{(i)}\right)\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)}{(T-t_{i})^{m/2}}\Biggr{)}\|\nabla f\|_{\infty}.\end{split}$ (26)
for some $Q>0$ and $s_{i}=t_{i}-t_{i-1}$.
Hereinafter, we refer to the first term of the right-hand side of (26)
@@
Let $\hat{Q}_{(s)}$ be an $m$-similar operator. From Theorem 3 of [13], there exists $Q\in\mathbb{N}$ such that
$\begin{split}\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{1}\left(s^{{(m+1)}/{2}}\|\nabla f\|_{\infty}+\sum_{q=m+1}^{K}s^{{q}/{2}}\|f\|_{V,q}\right)\end{split}$ (27)
@@
$\begin{split}\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{3}s^{1/2}\|\nabla f\|_{\infty}\end{split}$ (28)
hold for $s>0$, where $\|\cdot\|_{V,q}$ is a semi-norm on $C_{b}^{\infty}\left(\mathbb{R}^{N};\mathbb{R}\right)$ defined as
@@
We also have the following estimation from Theorem 2 of [12].
$\begin{split}\left\|P^{X}_{T-t}f\right\|_{V,q}\leq c_{2}\frac{\|\nabla f\|_{\infty}}{\left(T-t\right)^{(q-1)/2}}.\end{split}$ (30)
Then, setting $\nu^{(i)}=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right)$ as in (11), we decompose the total error as follows. Note that the operator $P^{X}_{T}$ and the measure $\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$, and its intermediate measures $\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$, $\nu^{(i)}$ depend on the initial value $x_{0}$.
$\begin{split}&\sup_{x_{0}\in\mathbb{R}^{N}}\left|P^{X}_{T}f(x)-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right|\\ &\qquad\qquad\quad\leq\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}P_{T-t_{1}}f\right\|_{\infty}\\ &\qquad\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\qquad\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right\|_{\infty}.\end{split}$ (31)
The first and second terms of right-hand side of (31), which correspond to the discretisation error, can be estimated as follows.
@@
Then, using (27), (28), (30), we have
$\begin{split}&\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}\left(P_{T-t_{1}}\right)\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\left\|P^{X}_{T-t_{i}}f\right\|_{V,q}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\frac{\left\|\nabla f\right\|_{\infty}}{\left(T-t_{i}\right)^{{q}/{2}}}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq C_{1}(T)\left({s_{n}}^{1/2}+\sum_{i=1}^{n-1}\sum_{q=m+1}^{K}\frac{{s_{i}}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\|\nabla f\|_{\infty}\end{split}$

### v2

@@ 2.3. N–V method
We introduce the N–V method from [24], which is a second-order discretisation scheme. Hereafter, we denote by $(G_{1}\circ G_{2})g(x)$ the composition $(G_{1}(G_{2}g))(x)$ of $G_{1}$ and $G_{2}$, where $G_{1},G_{2}\colon C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})\rightarrow C_{\rm{Lip}}(\mathbb{R}^{N};\mathbb{R})$.
@@
Let $\eta_{1},\eta_{2},\dots,\eta_{d}$ be independent standard normal random variables and define
@@
Then $Q^{(\rm{NV})}_{(s)}$ are 5-similar operators and the N–V method $Q^{(\rm{NV})}_{(T/n)}\circ Q^{(\rm{NV})}_{(T/n)}\circ\cdots\circ Q^{(\rm{NV})}_{(T/n)}$ is a second-order discretisation method.
Replacing the independent standard normal random variables $\left\{\eta_{i}\right\}_{i=1}^{d}$ in the definition above with the independent discrete random variables $\left\{\hat{\eta}_{i}^{(\text{$5$th})}\right\}_{i=1}^{d}$, we define the discrete N–V method as in [22]. Each $\hat{\eta}_{i}^{(\text{$5$th})}$ is a discrete random variable on $(\hat{\Omega},\hat{\mathcal{F}},\hat{P})$ defined by
$\hat{P}\left(\hat{\eta}_{i}^{(\text{$5$th})}=\pm\sqrt{3}\right)=\frac{1}{6},\;\;\hat{P}\left(\hat{\eta}_{i}^{(\text{$5$th})}=0\right)=\frac{2}{3}.$ (23)
We note that the moments of $\hat{\eta}_{i}^{(\text{$5$th})}$ up to the $5$th order match those of the standard normal distribution.
As in subsection 1.2.2, we denote a probability measure on $\mathbb{R}^{N}$ induced by $X^{({\rm Alg},\Delta)}\left(t_{i}\right)$ as $\mu_{t_{i}}^{({\rm Alg},\Delta)}=\sum_{j=1}^{m_{i}}w_{j}^{(i)}\delta_{p_{j}^{(i)}}$, where $\left\{\left(p_{j}^{(i)},w_{j}^{(i)}\right)\right\}_{j=1}^{m_{i}}$ is the set of pairs consisting of points $p_{j}^{(i)}$ in $\mathbb{R}^{N}$ and weights $w_{j}^{(i)}$ assigned to them. Let $\{\hat{\eta}^{(5{\rm th})}_{i,l}\}_{\begin{subarray}{c}i=1,\dots,n\\ l=1,\dots,d\end{subarray}}$ be a set of independent and identically distributed discrete random variables following the same distribution as $\hat{\eta}_{i}^{(5{\rm th})}$. Let $\phi$ be a function from {±$\{\pm\sqrt{3},0\}^{d}$ to $[0,1]$ defined by
$\phi(\eta)=\left(\frac{1}{6}\right)^{\operatorname{card}\{i\mid\eta_{i}=\pm\sqrt{3}\}}\left(\frac{2}{3}\right)^{\operatorname{card}\{i\mid\eta_{i}=0\}}$
for $\eta=\left(\eta_{1},\dots,\eta_{d}\right)\in\{\pm\sqrt{3},0\}^{d}$. We also define functions $y^{(+)}$ and $y^{(-)}$ from $\mathbb{R}^{N}\times\{\pm\sqrt{3},0\}^{d}\times\mathbb{R}_{>0}$ to $\mathbb{R}^{N}$ as
$\begin{split}y^{(+)}(x,\eta,s)&=\exp(sV_{0})\exp(\sqrt{s}\eta_{1}V_{1})\cdots\exp(\sqrt{s}\eta_{d}V_{d})(x),\\ y^{(-)}(x,\eta,s)&=\exp(\sqrt{s}\eta_{d}V_{d})\cdots\exp(\sqrt{s}\eta_{1}V_{1})\exp(sV_{0})(x).\end{split}$
Then, the operator $\hat{Q}_{(s_{i})}^{({\rm NV},\Delta)}$ between discrete measures is defined as
@@
In this way, we construct a concrete example of $\hat{Q}^{\left({\rm Alg},\Delta\right)}_{(s_{i})}$ used in this paper, as discussed in subsection 1.2.2. We then apply the recombination algorithm described in subsection 1.3.2 to this operator and obtain $\mu_{T}^{\left({\rm NV},\Delta,{\rm Rec}\right)}$.
We similarly define $\hat{Q}_{(s_{i})}(\mu)$ for an $m$-moment similar operator $Q_{(s)}$, using discrete random variables $\hat{\eta}^{(m\text{th})}$ whose moments up to the $m$th order match those of the standard normal random variables. In this paper, we consider a $p$th order discretisation algorithm defined by a $p$-moment similar operator and its associated recombination measure.
@@
In [24], the authors propose implementing the N–V method using Bernoulli random variables in addition to the standard normal random variables $\eta_{1},\eta_{2},\dots,\eta_{d}$ in order to halve the number of ODEs to be solved. In contrast, the present paper employs the N–V method without utilising Bernoulli random variables instead calculating the expectation as shown in equation (22). This approach is motivated by the need to minimise the number of support points in the discrete N–V method so as to regulate the overall computational cost, as noted in Remark 3.
@@
The N–V method offers the following advantages over other high-order discretisation methods, such as those in [20, 26]. As stated in the original paper [24], it is often more efficiently implementable because analytical solutions to the ODEs are available in certain cases, such as the Heston model [10]. Furthermore, the Feller condition [5] guarantees the positivity of processes discretised by the N–V method.
Another advantage is that, when implementing high-order discretisation methods using discrete random variables, the N–V method is economical in the sense that only $d$ random numbers are required to approximate an SDE driven by $d$-dimensional Brownian motion. The numbers of random variables involved and the cardinalities of the supports of the resulting measures are summarised in Table 1.
@@
Table 1. Number of random variables and support sizes for $\hat{X}^{(\text{method})}_{s,x}$, with $\sharp\Delta=n$.
@@ 2.4. High-order recombination: approximation theorem
In this subsection, we introduce the approximation theorem established in [17]
@@
The vector fields $V_{0},V_{1},\dots,V_{d}$ are said to satisfy the uniform Hörmandar (UH) condition if there exists $\ell\in\mathbb{N}$ such that
@@
Theorem 4 ([17], Theorem 19).
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th-order discretisation, and let $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let
$\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},\left\{U_{k}^{(i)}\right\}_{k=1}^{l_{i}}\right\}_{i=0}^{\sharp\Delta}$
be a set of recombination measures and patch divisions of their supports. Then the following inequality holds:
$\begin{split}&\left\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}(f)\right\|_{\infty}\\ &\qquad\leq\Bigg{(}C_{1}(T)\Bigg{(}{s_{\sharp\Delta}}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{j=m}^{m+1}\frac{{s_{i}}^{(j+1)/2}}{(T-t_{i})^{j/2}}\Bigg{)}\\ &\qquad\qquad\qquad\quad+C_{2}(T)\sum_{i=2}^{\sharp\Delta-1}\frac{{u_{i}}^{m+1}}{(T-t_{i})^{m\ell}}\Bigg{)}\|f\|_{\rm{Lip}}\end{split}$ (25)
where $u_{i}=\max\left.\left\{\operatorname{rad}\left(U_{k}^{(i)}\right)\;\right|\;k=1,\dots,l_{i}\right\}$ and $s_{i}=t_{i}-t_{i-1}$.
In the following, we present an example of a condition that satisfies the assumptions of the above theorem and can be used to determine the patch division.
Example 1 (LL patch condtion [17], Examples 21).
@@
$t_{i}=T\left(1-\left(1-\frac{j}{n}\right)^{\gamma}\right)$ (26)
@@
$u_{i}=\left(\frac{{s_{i}}^{m+1}}{(T-t_{i})^{m(1-\ell)}}\right)^{1/(2(m+1))},$ (27)
@@
In other words, the recombination does not compromise the order of the discretisation error.
We refer to the condition introduced above by Lyons–Litterer as the Lyons–Litterer patch condition, or simply the LL patch condition.
@@ 3. Refined error estimation and patch division algorithms
In this section, we present the main results of this paper. First, we refine Theorem 4 and Example 1 by extending the argument in [17], leading to Theorem 5 and Example 2, which gives a sufficient condition such that the patch division satisfies the required bound on the recombination error.
@@ 3.1. Refined error estimation
Taking the weight of each patch into account yields Theorem 5 below, which refines Theorem 4 by incorporating not only the radii of the patches but also their weights.
@@
Assume that the vector fields $V_{0},V_{1},\dots,V_{d}$ satisfy the UH condition. Let Alg be an $m$th-order discretisation, and let $\Delta$ be a partition of $[0,T]$ as $0=t_{0}<t_{1}<\dots<t_{\sharp\Delta}=T$. Let
$\left\{\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)},\left\{U_{k}^{(i)}\right\}_{k=1}^{l_{i}}\right\}_{i=0}^{\sharp\Delta}$
be a collection of recombined measures and patch divisions of their supports. Then, for some $Q>0$ and $s_{i}=t_{i}-t_{i-1}$, the following inequality holds:
$\begin{split}&\left\|P^{X}_{T}f-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(f\right)\right\|_{\infty}\\ &\leq\Biggl{(}C_{1}(T)\left({s_{\sharp\Delta}}^{1/2}+\sum_{i=1}^{\sharp\Delta-1}\sum_{q=m+1}^{K}\frac{{s_{i}}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\\ &\qquad\qquad\qquad+C_{2}(T)\sum_{i=1}^{\sharp\Delta-1}\frac{\sum_{k=1}^{l_{i}}\left(\operatorname{rad}\left(U_{k}^{(i)}\right)\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)}{(T-t_{i})^{m/2}}\Biggr{)}\|\nabla f\|_{\infty}.\end{split}$ (29)
Hereafter, we refer to the first term of the right-hand side of (29)
@@
Let $\hat{Q}_{(s)}$ be an $m$-similar operator. According to Theorem 3 of [13] there exists $Q\in\mathbb{N}$ such that
$\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{1}\left(s^{{(m+1)}/{2}}\|\nabla f\|_{\infty}+\sum_{q=m+1}^{K}s^{{q}/{2}}\|f\|_{V,q}\right)$ (30)
@@
$\left\|P^{X}_{s}f-\hat{Q}_{(s)}f\right\|_{\infty}\leq c_{3}s^{1/2}\|\nabla f\|_{\infty}$ (31)
hold for $s>0$, where $\|\cdot\|_{V,q}$ denotes a semi-norm on $C_{b}^{\infty}\left(\mathbb{R}^{N};\mathbb{R}\right)$ defined by
@@
Furthermore Theorem 2 of [12] yields the following estimate:
$\left\|P^{X}_{T-t}f\right\|_{V,q}\leq c_{2}\frac{\|\nabla f\|_{\infty}}{\left(T-t\right)^{(q-1)/2}}.$ (33)
Define $\nu^{(i)}=\hat{Q}_{(s_{i})}^{(\rm{Alg},\Delta)}\left(\mu_{t_{i-1}}^{(\rm{Alg},\Delta,\rm{Rec})}\right)$ as in (12). The total error can be decomposed as follows. Note that the operator $P^{X}_{T}$, the measure $\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$, and the intermediate measures $\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}$ and $\nu^{(i)}$ depend on the initial value $x_{0}$.
$\begin{split}&\sup_{x_{0}\in\mathbb{R}^{N}}\left|P^{X}_{T}f(x_{0})-\mu_{T}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}(f)\right|\\ &\qquad\qquad\leq\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}P_{T-t_{1}}f\right\|_{\infty}\\ &\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\qquad\qquad\quad+\sum_{i=1}^{n-1}\left\|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right\|_{\infty}.\end{split}$ (34)
The first and second terms of the right-hand side of (34), which correspond to the discretisation error, can be estimated as follows.
@@
Then, (30), (31), (33), yield
$\begin{split}&\left\|P^{X}_{T}f-\hat{Q}_{(s_{1})}\left(P_{T-t_{1}}\right)\right\|_{\infty}+\sum_{i=1}^{n-1}\left\|\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)-\nu^{(i+1)}\left(P^{X}_{T-t_{i+1}}f\right)\right\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\left\|P^{X}_{T-t_{i}}f\right\|_{V,q}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq\sum_{i=1}^{n-1}c_{1}\left({s_{i}}^{{(m+1)}/{2}}\left\|\nabla P^{X}_{T-t_{i}}f\right\|_{\infty}+\sum_{q=m+1}^{K}{s_{i}}^{{q}/{2}}\frac{\left\|\nabla f\right\|_{\infty}}{\left(T-t_{i}\right)^{{q}/{2}}}\right)+c_{3}{s_{n}}^{{1}/{2}}\|\nabla f\|_{\infty}\\ &\;\;\leq C_{1}(T)\left({s_{n}}^{1/2}+\sum_{i=1}^{n-1}\sum_{q=m+1}^{K}\frac{{s_{i}}^{q/2}}{(T-t_{i})^{(q-1)/2}}\right)\|\nabla f\|_{\infty}.\end{split}$